# Modular reduction of complex representations of finite reductive groups

### v1

@@ 3.3. The dual basis and the pseudo-dual basis
Let $\{f_{w}^{*}\}_{w\in W}$ be the dual basis to $\{f_{w}\}_{w\in W}$ under the perfect pairing $\langle$,$\rangle$ , ⟩.
@@
Of note for the purposes of studying Lusztig’s conjectures is the fact that in all of the cases for which Lusztig computed the values $M_{w}$ in [L], the equation (12) holds. This discrepancy between $f_{w}^{*}$ and $f^{w}$ in general type provides an explanation for the failure of [L, Conjecture 2.3ii)] for our elements $M_{w}$ as defined in Definition 4.1; this is the subject of our Section 5.2 to follow. That being said, Lemma 3.4 proves useful in computing $\{f_{w}^{*}\}_{w\in W}$ in terms of $\{f^{w}\}_{w\in W}$ in practice, as one can apply the Gram-Schmidt algorithm to the pseudo-dual basis $\{f^{y}\}_{y\in W}$ to compute $f_{w}^{*}$.
3.4. Conclusion in Type $A$
@@
where $w$ is some element of the unique two-sided cell for which $[\mathrm{gr}_{w}(N):\chi]\neq 0$. Since we are in Type A, $\mathrm{gr}_{w}(N)$ breaks into a direct sum of left cell representations, each of which is isomorphic to $\chi$ as a $\mathbb{C}[W]$-module; further, each has the form $\mathbb{C}[W]f_{y}$ for some unique involution $y\sim_{\mathrm{L}}w$. Since $\tilde{\phi}$ intertwines the $W$-action, we see that
@@ 4.1. Definition of the elements $M_{w}$
For any $w\in W$, let $\delta(w)$ be the degree of the Kazhdan-Lusztig polynomial $P_{1,w}$. It was shown in [Jos] that for each left cell of $W$, the function $w\mapsto\ell(w)-2\delta(w)$ reaches its minimum at a unique element of this cell, which is called the Duflo involution. It is shown in [LCellsII] that $\ell(w)-2\delta(w)=a(w)$ for Lusztig’s $a$-function on the Weyl group introduced in [LCellsI], which provides an alternate characterization of Duflo involutions.
@@
For any $w\in\mathcal{J}$, there exists some subset $\Lambda_{w}\subset\Lambda$ independent of $q$ and some coefficients $r_{\lambda}\in\mathbb{Z}$ such that
(16) $M_{w}=\sum_{\lambda\in\Lambda_{w}}r_{\lambda}V_{(q-1)\omega_{I(w)}+\lambda}.$
@@
Let $w\in\mathcal{J}$, and let $d$ be the Duflo involution in the same left cell as $w$. First, note that $[q]^{*}f_{w_{0}d}$ is a linear combination of monomials from the $W$-orbit of $\exp(q\omega_{I(w)})$, since recall that $\overline{I(w_{0}d)}=I(d)=I(w)$. The pairing of such a monomial $\exp(\nu)$ with another monomial $\exp(\mu)$ appearing with nonzero coefficient in $f_{w_{0}w}^{*}$ equals $V_{y(\nu+\mu)-\rho}=V_{(q-1)\omega_{I(w)}+\lambda}$ for some $\lambda$, where $y\in W$ is chosen such that $y(\nu+\mu)$ is dominant. We can write
$[q]^{*}f_{w_{0}d}$ $=\sum_{\nu\in W\cdot q\omega_{I(w)}}a_{\nu}\exp(\nu),$
@@
Then equation (16) holds, where $\Lambda_{w}$ is the set of all $\lambda$ for which $\mu+\nu$ and $(q-1)\omega_{I(w)}+\rho+\lambda$ lie in the same orbit, where $\mu\in W\cdot q\omega_{I(w)}$, $\nu\in\Lambda_{w}^{\prime}$, and for every such $\lambda$,
@@ 4.2. Cell modules and the asymptotic Hecke algebra
The basis $\{C_{w}\}_{w\in W}$ for $\mathbb{Z}[W]$ we consider in the present paper comes from the Kazhdan-Lusztig basis, which we still refer to as $C_{w}$, of the Hecke algebra $\mathcal{H}$ over $\mathbb{Z}[v,v^{-1}]$. We write $h_{x,y,z}\in\mathbb{Z}[v,v^{-1}]$ for the structure constants of multiplication in $\mathcal{H}$ with respect to the basis $\{C_{w}\}_{w\in W}$. Recalling again Lusztig’s $a$-function defined in [LCellsI] (which is constant on two-sided cells), we define for any $x,y,z\in W$ the integer $\gamma_{x,y,z}\in\mathbb{Z}$ to be the constant term of $v^{a(z)}h_{x,y,z^{-1}}\in\mathbb{Z}[v]$.
@@
One can show that there is a natural correspondence between irreducible modules over $\mathbb{C}[W]$, $\mathcal{H}$, and $J$, cf. for example the explanation given in [WheresCurtis, Sections 8, 9]. The following result appears in [LCellsII].
Proposition 4.3 ([LCellsII]).
@@
In the formulation of [L, Conjecture 2.3], Lusztig makes use of a virtual representation $R_{\alpha_{w}}$ defined in terms of the asymptotic Hecke algebra for any $w\in\mathcal{J}$. To recall it, we first recall the coefficients $c_{w,\chi}$ appearing in loc. cit.
@@
    For any irreducible representation $\chi$ of $W$ occuring with nonzero multiplicity in $\mathbb{Z}[W]_{c}$, let $c_{w,\chi}$ be the trace of $t_{w}$ in the corresponding representation of $J_{c}$.
@@
    $R_{\alpha_{w}}$ $=\frac{1}{|W|}\sum_{\begin{subarray}{c}\chi\in\mathrm{Irr}{W}\\ y\in W\end{subarray}}c_{w,\chi}\mathrm{tr}(y,\chi)R_{y},$
@@
    where $R_{y}$ is the unipotent Deligne-Lusztig character associated to $y$ as defined in [DL, Section 1.5].
@@ 4.3. Proof of Lusztig’s main conjecture
Now note that expanding both sides of (23) linearly gives the same result when $V_{\chi}$ is replaced by any linear combination of almost characters. Recall that for any $w\in W$, $R_{w}$ lies in the span of almost characters $V_{\chi}$; further, the value of any irreducible unipotent character $\chi$ on any semisimple conjugacy class is determined by $(\chi,R_{w})$ across all $w\in W$ cf. [Car, p. 383]. Since the Brauer reduction map vanishes on any class function on $G(\mathbb{F}_{q})$ supported away from semisimple elements, this implies that Corollary 4.7 is true for any irreducible unipotent representation. Thus we arrive at the following resolution to Conjecture 1.2.
@@ 5.1. Bounding weights in the Weyl module expansion
The second part of [L, Conjecture 2.3] asserts that the $V_{\lambda}$ which occur with nonzero coefficient in $M_{w}$ (as in Lemma 4.2) are such that $\lambda$ is “very close” to $\exp((q-1)\omega_{I(w)})$. In this section, we prove this part of the conjecture and provide an explicit bound.
@@
For any $w\in W$, $M_{w_{0}w}$ can be expressed as a linear combination of terms of the form $\mathrm{ch}(V_{\lambda})$ where $\lambda=(q-1)\omega_{\overline{I(w)}}+\mu$, where $\mu$ is in the $W$-orbit of some dominant weight between 00 and $2(h(w)+1)\rho$.
@@
By writing out $\langle f_{w}^{*},(f^{y})^{*}\rangle$ as in (29) and then expressing it as a linear combination of weights $\exp(\mu)$, those which can appear with nonzero coefficient are all in the $W$-orbit of some dominant weight between 00 and $2(h(w)-h(y))\rho$. This is because there are at most $h(w)-h(y)$ terms in the expression, and each term is itself a linear combination of weights in the $W$-orbit of a dominant weight between 00 and $2\rho$. In particular, the entire right-hand side of (31) is a linear combination of $\exp(\mu)$ for $\mu$ in the orbit of some dominant weight between 00 and $2h(w)\rho$.
Now note that $\langle[q]^{*}f_{d},f^{w}\rangle$ is a linear combination of $\exp(\mu)$ for $\mu$ in the $W$-orbit of weights of the form $q\omega_{\overline{I(w)}}+y\omega_{I(w)}$ for some $y\in W$. Adding the right-hand side of (31) to this term and writing $q\omega_{\overline{I(w)}}=(q-1)\omega_{\overline{I(w)}}+\omega_{\overline{I(w)}}$ we get that $\langle[q]^{*}f_{d},f_{w}^{*}\rangle$ can be written as a linear combination of $\exp(\mu)$ for
(32) $\mu=z((q-1)\omega_{\overline{I(w)}}+\mu^{\prime})$
for some $z\in W$, where $\mu^{\prime}$ is in the $W$-orbit of some dominant weight between 00 and $(2h(w)+1)\rho$. For $q>2h(w)+1$, it is the only dominant element in its $W$-orbit, and so when we express it as a linear combination of characters $V_{\lambda}$, we must have that
(33) $\lambda+\rho=(q-1)\omega_{\overline{I(w)}}+\mu^{\prime},$
where again $\mu^{\prime}$ is in the $W$-orbit of some dominant weight between 00 and $(2h(w)+1)\rho$. This means $\lambda-(q-1)\omega_{\overline{I(w)}}$ is in the $W$-orbit of some dominant weight between 00 and $2(h(w)+1)\rho$. ∎
@@
For any $w\in W$, $M_{w}$ can be expressed as a linear combination of terms of the form $\mathrm{ch}(V_{\lambda})$, where $\lambda-(q-1)\omega_{\overline{I}}$ lies in the $W$-orbit of some dominant weight less than or equal to $k\rho$, where $k=2(\rho,\rho)+2$.
@@ 5.2. Conjectured symmetries for dimension polynomials
For the $M_{w}$ which were already computed in [L] in types $A_{1}$, $A_{2}$, $B_{2}$, $G_{2}$, $A_{3}$, this conjecture holds. Despite our $M_{w}$ agreeing with those computations, these elements fail to satisfy this symmetry in Type $A_{4}$.
@@
In Table 6, we list all of the values $P_{w}(q)$ for $w\in W$, and clearly none of them are of the form ±$\pm t^{\nu}P_{w}(1/q)$ when $P_{w}(q)$ is either of the two polynomials above. ∎
@@
The evaluation at $1\in T/W$ is a surjective homomorphism $N\twoheadrightarrow N\otimes_{A}\mathbb{C}\cong H^{\bullet}(G/B)$. The endomorphism $[q]^{*}$ induces the endomorphism $[q]^{*}_{1}$ of $H^{\bullet}(G/B)$ that acts by $q^{i}$ on $H^{2i}(G/B)$. Since $M_{w}$ was a matrix entry of $[q]^{*}$ in the basis $\{f_{y},\ y\in W\}$, the dimension of $M_{w}$ is a matrix entry of $[q]^{*}_{1}$ in the basis $\{(f_{y})|_{1},\ y\in W\}\subset H^{\bullet}(G/B)$. It is known that the $W$-representations in a cell $c$ appear in the cohomology $H^{2i}(G/B)$ only for $i\geq a(c)$. Since $a(c^{\prime})\leq a(c)$ for $c^{\prime}\leq c$, the desired divisibility follows. ∎
@@ 5.3. Positivity
There exist examples of $M_{w}$ which do not lie in $R_{p}^{+}G(\mathbb{F}_{q})$. For example, in Type $B_{3}$ for $w=s_{1}$, and in Type $A_{4}$ for $w=s_{2}s_{3}s_{2}$.
@@
We used the software [L] to compute the following expansions for $M_{w}$ in terms of the irreducible representations $L_{\lambda}$.
@@
For all the types for which $M_{w}$ was computed in [L] (Types $A_{1}$, $A_{2}$, $B_{2}$, $G_{2}$, $A_{3}$), positivity is satisfied. For example, in Table 2 for Type $A_{3}$, we explicitly write each $M_{w}$ in terms of the basis $L_{\lambda}$ of irreducibles, exhibiting positivity in this case. We used the software [dynkin] to produce this table.

### v2

@@ 3.3. The dual basis and the pseudo-dual basis
Let $\{f_{w}^{*}\}_{w\in W}$ be the dual basis to $\{f_{w}\}_{w\in W}$ under the perfect pairing $\langle,\rangle$.
@@
Of note for the purposes of studying Lusztig’s conjectures is the fact that in all of the cases for which Lusztig computed the values $M_{w}$ in [undefp], the equation (12) holds. This discrepancy between $f_{w}^{*}$ and $f^{w}$ in general type provides an explanation for the failure of [undefp, Conjecture 2.3ii)] for our elements $M_{w}$ as defined in Definition 4.1; this is the subject of our Section 5.2 to follow. That being said, Lemma 3.4 proves useful in computing $\{f_{w}^{*}\}_{w\in W}$ in terms of $\{f^{w}\}_{w\in W}$ in practice, as one can apply the Gram-Schmidt algorithm to the pseudo-dual basis $\{f^{y}\}_{y\in W}$ to compute $f_{w}^{*}$.
3.4. Conclusion in Type AA
@@
where $w$ is some element of the unique two-sided cell for which $\mathrm{gr}_{w}(N):\chi]\neq 0$. Since we are in Type A, $\mathrm{gr}_{w}(N)$ breaks into a direct sum of left cell representations, each of which is isomorphic to $\chi$ as a $\mathbb{C}[W]$-module; further, each has the form $\mathbb{C}[W]f_{y}$ for some unique involution $y\sim_{\mathrm{L}}w$. Since $\tilde{\phi}$ intertwines the $W$-action, we see that
@@ 4.1. Definition of the elements $M_{w}$
For any $w\in W$, let $\delta(w)$ be the degree of the Kazhdan-Lusztig polynomial $P_{1,w}$. It was shown in [undefj] that for each left cell of $W$, the function $w\mapsto\ell(w)-2\delta(w)$ reaches its minimum at a unique element of this cell, which is called the Duflo involution. It is shown in [undefr] that $\ell(w)-2\delta(w)=c(w)$ for $c(w)$ in the present paper denoting the value on $w$ of Lusztig’s $a$-function on the Weyl group introduced in [undefq], which provides an alternate characterization of Duflo involutions.
@@
For any $w\in\mathcal{J}$, there exists some subset $\Lambda_{w}\subset X^{*}(T)$ independent of $q$ and some coefficients $r_{\lambda}\in\mathbb{Z}$ such that
(16) $M_{w}=\sum_{\lambda\in\Lambda_{w}}r_{\lambda}V_{(q-1)\varpi_{I(w)}+\lambda}.$
@@
Let $w\in\mathcal{J}$, and let $d$ be the Duflo involution in the same left cell as $w$. First, note that $q]^{*}f_{w_{0}d}$ is a linear combination of monomials from the $W$-orbit of $\exp(q\varpi_{I(w)})$, since recall that $\overline{I(w_{0}d)}=I(d)=I(w)$. The pairing of such a monomial $\exp(\nu)$ with another monomial $\exp(\mu)$ appearing with nonzero coefficient in $f_{w_{0}w}^{*}$ equals $V_{y(\nu+\mu)-\rho}=V_{(q-1)\varpi_{I(w)}+\lambda}$ for some $\lambda$, where $y\in W$ is chosen such that $y(\nu+\mu)$ is dominant. We can write
$[q]^{*}f_{w_{0}d}$ $=\sum_{\nu\in W\cdot q\varpi_{I(w)}}a_{\nu}\exp(\nu),$
@@
Then equation (16) holds, where $\Lambda_{w}$ is the set of all $\lambda$ for which $\mu+\nu$ and $(q-1)\varpi_{I(w)}+\rho+\lambda$ lie in the same orbit, where $\nu\in W\cdot q\varpi_{I(w)}$, $\mu\in\Lambda_{w}^{\prime}$, and for every such $\lambda$,
@@ 4.2. Cell modules and the asymptotic Hecke algebra
The basis $\{C_{w}\}_{w\in W}$ for $\mathbb{Z}[W]$ we consider in the present paper comes from the Kazhdan-Lusztig basis, which we still refer to as $C_{w}$, of the Hecke algebra $\mathcal{H}$ over $\mathbb{Z}[v,v^{-1}]$. We write $h_{x,y,z}\in\mathbb{Z}[v,v^{-1}]$ for the structure constants of multiplication in $\mathcal{H}$ with respect to the basis $\{C_{w}\}_{w\in W}$. Recalling again Lusztig’s $a$-function defined in [undefq] (which is constant on two-sided cells), we define for any $x,y,z\in W$ the integer $\gamma_{x,y,z}\in\mathbb{Z}$ to be the constant term of $v^{a(z)}h_{x,y,z^{-1}}\in\mathbb{Z}[v]$.
@@
One can show that there is a natural correspondence between irreducible modules over $\mathbb{C}[W]$, $\mathcal{H}$, and $J$, cf. for example the explanation given in [undefg, Sections 8, 9]. The following result appears in [undefr].
Proposition 4.3 ([undefr]).
@@
In the formulation of [undefp, Conjecture 2.3], Lusztig makes use of a virtual representation $R_{\alpha_{w}}$ defined in terms of the asymptotic Hecke algebra for any $w\in\mathcal{J}$. To recall it, we first recall the coefficients $c_{w,\chi}$ appearing in loc. cit.
@@
    For any irreducible representation $\chi$ of $W$ occurring with nonzero multiplicity in $\mathbb{Z}[W]_{c}$, let $c_{w,\chi}$ be the trace of $t_{w}$ in the corresponding representation of $J_{c}$.
@@
    $R_{\alpha_{w}}$ $=\frac{1}{|W|}\sum_{\begin{subarray}{c}\chi\in\mathrm{Irr}(W)\\ y\in W\end{subarray}}c_{w,\chi}\mathrm{tr}(y,\chi)R_{y},$
@@
    where $R_{y}$ is the unipotent Deligne-Lusztig character associated to $y$ as defined in [undefh, Section 1.5].
@@ 4.3. Proof of Lusztig’s main conjecture
Now note that expanding both sides of (23) linearly gives the same result when $V_{\chi}$ is replaced by any linear combination of almost characters. Recall that for any $w\in W$, $R_{w}$ lies in the span of almost characters $V_{\chi}$; further, the value of any irreducible unipotent character $\chi$ on any semisimple conjugacy class is determined by $(\chi,R_{w})$ across all $w\in W$ cf. [undeff, p. 383]. Since the Brauer reduction map vanishes on any class function on $G(\mathbb{F}_{q})$ supported away from semisimple elements, this implies that Corollary 4.7 is true for any irreducible unipotent representation. Thus we arrive at the following resolution to Conjecture 1.1.
@@ 5.1. Bounding weights in the Weyl module expansion
The second part of [undefp, Conjecture 2.3] asserts that the $V_{\lambda}$ which occur with nonzero coefficient in $M_{w}$ (as in Lemma 4.2) are such that $\lambda$ is “very close” to $\exp((q-1)\varpi_{I(w)})$. In this section, we prove this part of the conjecture and provide an explicit bound.
@@
For any $w\in W$, $M_{w_{0}w}$ can be expressed as a linear combination of terms of the form $\mathrm{ch}(V_{\lambda})$ where $\lambda=(q-1)\varpi_{\overline{I(w)}}+\mu$, where $\mu$ is in the $W$-orbit of some dominant weight between 0 and $2(h(w)+1)\rho$.
@@
By writing out $\langle f_{w}^{*},(f^{y})^{*}\rangle$ as in (29) and then expressing it as a linear combination of weights $\exp(\mu)$, those which can appear with nonzero coefficient are all in the $W$-orbit of some dominant weight between 0 and $2(h(w)-h(y))\rho$. This is because there are at most h(w)h(y)h(w)-h(y) terms in the expression, and each term is itself a linear combination of weights in the $W$-orbit of a dominant weight between 0 and $2\rho$. In particular, the entire right-hand side of (31) is a linear combination of $\exp(\mu)$ for $\mu$ in the orbit of some dominant weight between 0 and $2h(w)\rho$.
Now note that $\langle[q]^{*}f_{d},f^{w}\rangle$ is a linear combination of $\exp(\mu)$ for $\mu$ in the $W$-orbit of weights of the form $q\varpi_{\overline{I(w)}}+y\varpi_{I(w)}$ for some $y\in W$. Adding the right-hand side of (31) to this term and writing $q\varpi_{\overline{I(w)}}=(q-1)\varpi_{\overline{I(w)}}+\varpi_{\overline{I(w)}}$ we get that $\langle[q]^{*}f_{d},f_{w}^{*}\rangle$ can be written as a linear combination of $\exp(\mu)$ for
(32) $\mu=z((q-1)\varpi_{\overline{I(w)}}+\mu^{\prime})$
for some $z\in W$, where $\mu^{\prime}$ is in the $W$-orbit of some dominant weight between 0 and $(2h(w)+1)\rho$. For $q>2h(w)+1$, it is the only dominant element in its $W$-orbit, and so when we express it as a linear combination of characters $V_{\lambda}$, we must have that
(33) $\lambda+\rho=(q-1)\varpi_{\overline{I(w)}}+\mu^{\prime},$
where again $\mu^{\prime}$ is in the $W$-orbit of some dominant weight between 0 and $(2h(w)+1)\rho$. This means $\lambda-(q-1)\varpi_{\overline{I(w)}}$ is in the $W$-orbit of some dominant weight between 0 and $2(h(w)+1)\rho$. ∎
@@
For any $w\in W$, $M_{w}$ can be expressed as a linear combination of terms of the form $\mathrm{ch}(V_{\lambda})$, where $\lambda-(q-1)\varpi_{\overline{I(w)}}$ lies in the $W$-orbit of some dominant weight less than or equal to $k\rho$, where $k=2(\rho,\rho)+2$.
@@ 5.2. Conjectured symmetries for dimension polynomials
For the $M_{w}$ which were already computed in [undefp] in types $A_{1}$, $A_{2}$, $B_{2}$, $G_{2}$, $A_{3}$, this conjecture holds. Despite our $M_{w}$ agreeing with those computations, these elements fail to satisfy this symmetry in Type $A_{4}$.
@@
In Table 6, we list all of the values $P_{w}(q)$ for $w\in W$, and clearly none of them are of the form $\pm t^{\nu}P_{w}(1/t)$ when $P_{w}(t)$ is either of the two polynomials above. ∎
@@
The evaluation at $1\in T/W$ is a surjective homomorphism $N\twoheadrightarrow N\otimes_{A}\mathbb{C}\cong H^{\bullet}(G/B)$. The endomorphism $q]^{*}$ induces the endomorphism $q]^{*}_{1}$ of $H^{\bullet}(G/B)$ that acts by $q^{i}$ on $H^{2i}(G/B)$. Since $M_{w}$ was a matrix entry of $q]^{*}$ in the basis $\{f_{y},\ y\in W\}$, the dimension of $M_{w}$ is a matrix entry of $q]^{*}_{1}$ in the basis $\{(f_{y})|_{1},\ y\in W\}\subset H^{\bullet}(G/B)$. It is known that the $W$-representations in a cell $c$ appear in the cohomology $H^{2i}(G/B)$ only for $i\geq a(c)$. Since $a(c^{\prime})\leq a(c)$ for $c^{\prime}\leq c$, the desired divisibility follows. ∎
@@ 5.3. Positivity
There exist examples of $M_{w}$ which do not lie in $\mathcal{R}_{p}^{+}G(\mathbb{F}_{q})$. For example, in Type $B_{3}$ for $w=s_{1}$, and in Type $A_{4}$ for $w=s_{2}s_{3}s_{2}$.
@@
We used the software [undefe] to compute the following expansions for $M_{w}$ in terms of the irreducible representations $L_{\lambda}$.
@@
For all the types for which $M_{w}$ was computed in [undefp] (Types $A_{1}$, $A_{2}$, $B_{2}$, $G_{2}$, $A_{3}$), positivity is satisfied. For example, in Table 2 for Type $A_{3}$, we explicitly write each $M_{w}$ in terms of the basis $L_{\lambda}$ of irreducibles, exhibiting positivity in this case. We used the software [undefe] to produce this table.